In 20 years, you are hoping to have saved $100,000 towards your child’s college education. If you are able to save $2,500 at the end of each year for the next 20 years, what rate of return must you earn on your investments in order to achieve your goal?

I want know How i can get the answer by using formula please?
Answer #1

use future value annuity formula

FV=100000

P*(((1+r)^n-1)/r)=100000

2500*(((1+r)^20-1)/r)=100000

(((1+r)^20-1)/r)=100000/2500

(((1+r)^20-1)/r)=40

use trial and error method to find r

after trying many values, we get r=6.77% (answer)

the above should be answer..